OpenStudy (rosales0228):

HEED HELP!!!! WILL GIVE MEDAL AND FAN!! Holly wants to save money for an emergency. Holly invests 1,500 in an account that pays in interest rate of 6.75% How many years will it take for the account to reach 7,300 round your answer to the nearest hundredth

OpenStudy (rosales0228):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Assuming you use the same formula A = P(1+r)^t, you'd have A = 7300 P = 1500 r = 0.0675 t = x

jimthompson5910 (jim_thompson5910):

A = P(1+r)^t 7300 = 1500(1+0.0675)^x solve for x

OpenStudy (rosales0228):

okay thats what i was doing wrong! i was doing everything the same besides 0.0675 i was doing 0.675

jimthompson5910 (jim_thompson5910):

0.675 = 67.5% I wish interest rates were that high

jimthompson5910 (jim_thompson5910):

then again, someone would have to be paying that, so nvm